# Distilling Diffusion Models to Efficient 3D LiDAR Scene Completion

### v1

@@ D.1 More ablation study of structural loss
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
ScoreLiDAR 0.342 0.399 0.452 0.437
w/o Point-wise loss 0.351 0.414 0.485 0.486
w/o Scene-wise loss 0.367 0.422 0.477 0.451
w/o Structural Loss 0.419 0.430 0.549 0.445
@@
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
$\lambda_{scene}=0.5,\lambda_{point}=0.01$ 0.342 0.399 0.452 0.437
$\lambda_{scene}=0.05,\lambda_{point}=0.01$ 0.354 0.409 0.494 0.457
$\lambda_{scene}=0.5,\lambda_{point}=0.1$ 0.358 0.419 0.539 0.474
@@ D.2 Ablation study of different key point number
CD $\downarrow$ JSD $\downarrow$ CD $\downarrow$ JSD $\downarrow$
$n=1/30$ 0.342 0.399 0.452 0.437
$n=1/60$ 0.346 0.409 0.452 0.471
Table 8: Ablation study of different key points number. The first row refers to the default configuration of the ScoreLiDAR. The metrics refer to the performance with refinement.
As mentioned in Sec. A.3, the optimal number of the key point is set to the $\frac{1}{30}$ of the total number of points in the ground truth. To investigate the impact of different numbers of key points on the completion performance of ScoreLiDAR, we decreased the number of key points for model training and evaluated the completion performance. As shown in Tab. 8, when the number of key points decreases, the performance of ScoreLiDAR declines. This is because an insufficient number of key points causes the point-wise loss to fail in effectively capturing the relative positional information between key points, preventing the student model from learning the local geometric structure, and thereby reducing the completion quality.
D.3 Ablation study of different sampling steps on KITTI-360
@@
Model CD $\downarrow$ JSD $\downarrow$ Time (s) $\downarrow$
LiDiff (50 steps) [24] 0.564 0.549 29.18
LiDiff (50 steps Refined) [24] 0.517 0.446 29.43
LiDiff (8 steps) [24] 0.619 0.471 5.46
LiDiff (8 steps Refined) [24] 0.550 0.462 5.77
ScoreLiDAR (8 Steps) 0.452 0.437 5.14
ScoreLiDAR (4 Steps) 0.482 0.461 3.16
ScoreLiDAR (2 Steps) 0.525 0.457 1.69
ScoreLiDAR (1 Steps) 0.750 0.478 1.03
@@ D.5 Visualization of key points
D.6 Experiments on scene occupancy
We calculate the Intersection-Over-Union (IoU) [31] to evaluate the occupancy of the completed scene compared with the ground truth scene. IoU represents the degree of overlap between the voxels in the completed scene and those in the ground truth scene. A higher IoU value indicates a higher completeness of the completed scene. During the evaluation, we considered three different voxel resolutions: $0.5m$, $0.2m$, and $0.1m$. The smaller the voxel resolution, the more fine-grained details are considered in the evaluation metrics, and vice versa.
Tab. 11 and Tab. 12 show the IoU of ScoreLiDAR and existing models. Under low voxel resolutions, ScoreLiDAR achieves comparable IoU values, meaning ScoreLiDAR generates dense and accurate point clouds. When the voxel resolutions become higher, the performance of ScoreLiDAR declines. This is because the existing method is mainly based on signed distance fields, which implement the scene completion using a voxel representation. ScoreLiDAR is point-level scene completion with the input of point clouds obtained from LiDAR scans, which works better at smaller voxel resolutions. The results in Tab. 11 and Tab. 12 do not align with our experimental results and user study findings. Therefore, these results are provided for reference only.
Appendix E Introduction on utilized methods
E.1 Variational score distillation
@@ E.2 MinkowskiEngine
Appendix F Ethical statement
@@
F.1 Notification to human subjects
@@
Appendix G Failure examples

### v2

@@ D.1 More ablation study of structural loss
CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$
ScoreLiDAR 0.342 0.399 23.26 0.452 0.437 23.02
w/o Point-wise loss 0.351 0.414 23.37 0.485 0.486 23.29
w/o Scene-wise loss 0.367 0.422 24.89 0.477 0.451 24.69
w/o Structural Loss 0.419 0.430 24.13 0.549 0.445 24.56
@@
CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$
$\lambda_{scene}=0.5,\lambda_{point}=0.01$ 0.342 0.399 23.26 0.452 0.437 23.02
$\lambda_{scene}=0.05,\lambda_{point}=0.01$ 0.354 0.409 23.40 0.494 0.457 23.21
$\lambda_{scene}=0.5,\lambda_{point}=0.1$ 0.358 0.419 23.27 0.539 0.474 23.09
@@ D.2 Ablation study of different key point number
CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$
$n=1/20$ 0.329 0.392 24.35 0.528 0.475 24.01
$n=1/30$ 0.342 0.399 23.26 0.452 0.437 23.02
$n=1/60$ 0.346 0.409 25.19 0.452 0.471 25.05
$n=1/70$ 0.428 0.454 25.60 0.466 0.479 25.44
Table 8: Ablation study of different key points number. The result of $n=1/30$ refers to the default configuration of the ScoreLiDAR. The metrics refer to the performance with refinement.
As mentioned in Sec. A.3, the optimal number of the key point is set to the $\frac{1}{30}$ of the total number of points in the ground truth. To investigate the impact of different numbers of key points on the completion performance of ScoreLiDAR, we decreased the number of key points for model training and evaluated the completion performance. As shown in Tab. 8, the final performance of ScoreLiDAR is positively correlated with the number of key points. When the number of key points decreases, the performance of ScoreLiDAR declines. This is because an insufficient number of key points causes the point-wise loss to fail in effectively capturing the relative positional information between key points, preventing the student model from learning the local geometric structure, and thereby reducing the completion quality. However, when the number of key points is too large, it can easily cause an out-of-memory issue and reduce training efficiency. Therefore, this paper sets $n=\frac{1}{30}$.
D.3 Ablation study of different key point selection method
We compare different key point selection methods including random selection and farthest selection with the proposed selection method based on curvature. The results in Tab. 9 show that using the proposed selection method based on curvature achieves the optimal performance than other selection methods.
Model CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$
Random selection 0.384 0.433 23.23
farthest selection 0.442 0.459 23.63
ScoreLiDAR 0.342 0.399 23.26
Table 9: Comparison of different selection methods on SemanticKITTI dataset. The performance of the proposed selection method is better than others.
D.4 Ablation study of different sampling steps on KITTI-360
@@
Model CD $\downarrow$ JSD $\downarrow$ EMD $\downarrow$ Time (s) $\downarrow$
LiDiff (50 steps) [28] 0.564 0.549 21.98 29.18
LiDiff (50 steps Refined) [28] 0.517 0.446 22.96 29.43
LiDiff (8 steps) [28] 0.619 0.471 24.85 5.46
LiDiff (8 steps Refined) [28] 0.550 0.462 25.49 5.77
ScoreLiDAR (8 Steps) 0.452 0.437 23.02 5.14
ScoreLiDAR (4 Steps) 0.482 0.461 23.76 3.16
ScoreLiDAR (2 Steps) 0.525 0.457 - 1.69
ScoreLiDAR (1 Steps) 0.750 0.478 - 1.03
@@ D.6 Visualization of key points
D.7 Experiments on semantic scene completion
The objective of this paper is to propose a foundational method for distillation acceleration applicable to various LiDAR scene completion diffusion models. Semantic scene completion is not within the scope of the fundamental experiments considered in this paper. In spite of this, to validate the generalizability of the proposed method, we use SemCity [15] as the teacher model to verify the effectiveness of ScoreLiDAR on semantic scene completion tasks. Because the pre-trained models and code for the metric computation of SemCity are not publicly available, we retrained SemCity based on the official implementation and reproduced the metric computation ourselves. Tab. 14 shows the results. In semantic scene completion tasks, the proposed ScoreLiDAR still shows better completion quality than that of the teacher model.
Model FID $\downarrow$ KID $\downarrow$
SemCity [15] 88.52 0.11
ScoreLiDAR 81.76 0.09
Table 14: Results of semantic scene completion. Our ScoreLiDAR shows better performance.
D.8 Experiments on scene occupancy
We calculate the Intersection-Over-Union (IoU) [35] to evaluate the occupancy of the completed scene compared with the ground truth scene. IoU represents the degree of overlap between the voxels in the completed scene and those in the ground truth scene. A higher IoU value indicates a higher completeness of the completed scene. During the evaluation, we considered three different voxel resolutions: $0.5m$, $0.2m$, and $0.1m$. The smaller the voxel resolution, the more fine-grained details are considered in the evaluation metrics, and vice versa. However, IoU is a voxel-based metric, and in some voxel-based LiDAR scene completion methods, it can serve as an accurate measure of completion quality. In contrast, ScoreLiDAR is a point cloud-based completion method, which differs from traditional voxel-based approaches. As a result, IoU may introduce bias when evaluating the completion quality. Therefore, here we provide IoU results solely as a relative reference.
Tab. 12 and Tab. 13 show the IoU of ScoreLiDAR and existing models. Under low voxel resolutions, ScoreLiDAR achieves comparable IoU values, meaning ScoreLiDAR generates dense and accurate point clouds. When the voxel resolutions become higher, the performance of ScoreLiDAR declines. As mentioned above, the existing method is mainly based on signed distance fields, which implement the scene completion using a voxel representation. ScoreLiDAR is point-level scene completion with the input of point clouds obtained from LiDAR scans, which works better at smaller voxel resolutions.
Appendix E Theoretical demonstration
The gradient of the student model is Eq. 25.
As proposed in ScoreSDE [36], the log likelihood $\nabla_{\mathcal{G}^{t}}\log q^{t}\left(\mathcal{G}^{t}\right)$ can be approximated by the predicted noise $\hat{\epsilon}$ with $\nabla_{\mathcal{G}^{t}}\log q^{t}\left(\mathcal{G}^{t}\right)\approx-\frac{\boldsymbol{\hat{\epsilon}}}{\sqrt{1-\bar{\alpha}^{t}}}$. Thus, the gradient in Eq. 25 can be written as
$\approx\mathbb{E}_{t,\epsilon}\left[-\frac{\boldsymbol{\epsilon}_{\phi}\left(\mathcal{G}^{t},\mathcal{P},t\right)}{\sqrt{1-\bar{\alpha}^{t}}}-\left(-\frac{\boldsymbol{\epsilon}_{\theta}\left(\mathcal{G}^{t},\mathcal{P},t\right)}{\sqrt{1-\bar{\alpha}^{t}}}\right)\right]\frac{\partial\mathcal{G}^{t}}{\partial\eta}$
$=\mathbb{E}_{t,\epsilon}\left[\frac{\boldsymbol{\epsilon}_{\theta}\left(\mathcal{G}^{t},\mathcal{P},t\right)}{\sqrt{1-\bar{\alpha}^{t}}}-\frac{\boldsymbol{\epsilon}_{\phi}\left(\mathcal{G}^{t},\mathcal{P},t\right)}{\sqrt{1-\bar{\alpha}^{t}}}\right]\frac{\partial\mathcal{G}^{t}}{\partial\eta}$
Here $\sqrt{1-\bar{\alpha}^{t}}$ can be ignored. Thus, the gradient of $G_{stu}$ can be approximated by
Appendix F Introduction on utilized methods
F.1 Variational score distillation
@@ F.2 MinkowskiEngine
Appendix G Ethical statement
@@
G.1 Notification to human subjects
@@
Appendix H Failure examples